I need to know how to find a line perpendicular to y=3x+6
Answer by mananth(16949)

1 y = 3 x + 6
Divide by 1
y = 3 x + 6
Compare this equation with y=mx+b
slope m = 3.00

The slope of a line perpendicular to the above line will be the negative reciprocal -0.33
m1*m2=-1
The slope of the required line will be - 1/3
You need to know the point the line passes through.
plug slope and the point in the equation y= mx+b
Find b
plug slope and b in y= mx+b
m.ananth@hotmail.ca